### v2

# Automorphism groups of affine varieties and their Lie algebras

Abstract.
This is a brief overview of a few selected chapters on automorphism groups of affine varieties. It includes some open problems.
Mathematics Subject Classification: 14J50, 14R20, 14L30, 14E07, 22F50.     Key words: affine variety, flexible variety, automorphism group, ind-group, Lie algebra, algebraic group action, one-parameter subgroup, Tits alternative, triangulation, linearization, amalgamation.
1. Introduction
We provide a brief overview of a few selected chapters on automorphism groups of affine varieties. It completes the existing literature on the subject, see e.g. [Arz23], [Fre17], [FuKr18], [Kal09], [KPZ17], [Kra89], [Kra96], [Kra17], [Lam24], [Miy24], [Sno89], [vdE00], [vdEKC21], [Wri17], etc.
Throughout the text, $X$ stands for an affine variety defined over an algebraically closed field ${\mathbb{K}}$ of characteristic zero, $\operatorname{Aut}(X)$ stands for the group of automorphisms of $X$, ${\mathbb{A}}^{n}$ stands for the affine $n$-space over ${\mathbb{K}}$ and $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$ resp. $\operatorname{{\mathbb{G}}_{\mathrm{m}}}$ stands for the additive resp. multiplicative group of the field ${\mathbb{K}}$.
If $\dim(X)=1$ then $\operatorname{Aut}(X)$ is an algebraic group. Moreover, any finite group can be realized as the full automorphism group of a smooth affine curve, see [Jel15, Proposition 7.2]. The properties of $\operatorname{Aut}(X)$ are ultimately related with the geometry of $X$. For instance, if $\operatorname{Aut}(X)$ is infinite, then $X$ is uniruled, i.e. covered by rational curves, see [Jel15, Theorem 1.1]. Already in the case of a smooth affine surface $X$, the group $\operatorname{Aut}(X)$ is often infinite-dimensional and has a rich algebraic structure, see the survey article [KPZ17]. The present survey is focused on various aspects of $\operatorname{Aut}(X)$ as an ind-group and on its Lie algebra. It contains a list of open problems partially borrowed in [FuKr18] and [KrZa24].
[…]
The group of regular automorphisms $\operatorname{Aut}(X)$ is an affine ind-group, see e.g. [FuKr18, Theorem 5.1.1]. A closed algebraic subset of $A_{i}$ defines an algebraic family of automorphisms of $X$; such families were studied in [Ram64], see also [FuKr18, Section 3.3] and [Pop14].
A subgroup $G\subseteq\operatorname{Aut}(X)$ that is a closed algebraic subset of some $A_{i}$ has a natural structure of an algebraic group. It will be called an algebraic subgroup of $\operatorname{Aut}(X)$. Since $\operatorname{Aut}(X)$ is an affine ind-group, any algebraic subgroup of $\operatorname{Aut}(X)$ is an affine algebraic group. A faithful action of a linear algebraic group $G$ on $X$ defines an isomorphism of $G$ with an algebraic subgroup of $\operatorname{Aut}(X)$.
[…]
Let $\mathfrak{G}=\operatorname{Aut}(X)$ and let $\operatorname{Vec}(X)$ stand for the Lie algebra of regular vector fields on $X$. Recall that $\operatorname{Vec}(X)$ is naturally isomorphic to the Lie algebra $\operatorname{Der}(\mathcal{O}(X))$ of derivations of the ${\mathbb{K}}$-algebra $\mathcal{O}(X)$ of regular functions on $X$. In turn, $\operatorname{Lie}(\operatorname{Aut}(X))$ can be naturally identified with a subalgebra of $\operatorname{Vec}(X)$, see [FuKr18, Propositions 3.2.4 and 7.2.4]. For $X={\mathbb{A}}^{n}$ this is the subalgebra $\operatorname{Vec}^{\mathrm{c}}({\mathbb{A}}^{n})\subseteq\operatorname{Vec}({\mathbb{A}}^{n})$ of all vector fields on ${\mathbb{A}}^{n}$ with constant divergence, see [Sha81, p. 191] and [FuKr18, Proposition 15.7.2]. For an algebraic subgroup $G\subseteq\operatorname{Aut}(X)$, $\operatorname{Lie}(G)$ is a finite-dimensional Lie subalgebra of $\operatorname{Vec}(X)$.
Let $\mathfrak{G}$ be an ind-group and $\mathfrak{H}\subseteq\mathfrak{G}$ be a closed ind-subgroup. Then $\operatorname{Lie}(\mathfrak{H})$ is a Lie subalgebra of $\operatorname{Lie}(\mathfrak{G})$. However, it can happen (contrary to [Sha81, Theorem 1]) that $\operatorname{Lie}(\mathfrak{H})=\operatorname{Lie}(\mathfrak{G})$ while $\mathfrak{H}$ is a proper subgroup of a connected ind-group $\mathfrak{G}$, see Example 9.5. In this respect, the ind-groups are more complex than the algebraic groups.
2. Lie algebras of vector fields
2.1. Algebraically generated subgroups of $\operatorname{Aut}(X)$ and Lie algebras
Let $G\subseteq\operatorname{Aut}(X)$ be a subgroup. One says that $G$ is algebraically generated if it is generated by a family $\{G_{i}\,|\,i\in I\}$ of connected algebraic subgroups of $\operatorname{Aut}(X)$. Thus, an algebraically generated subgroup is connected.
Throughout this subsection we assume that $G$ is algebraically generated. We associate to $G$ the Lie subalgebra
$L(G)=\langle\operatorname{Lie}(G_{i})\,|\,i\in I\rangle_{\tiny Lie}\ \subseteq\operatorname{Vec}(X)$
generated by the Lie subalgebras $\operatorname{Lie}(G_{i})$. In fact, $L(G)$ depends only on $G$ and not on the choice of generating subgroups $G_{i}$, see [KrZa24, Theorem 2.3.1(4)].
For any ind-group $\mathfrak{G}$ we also define the canonical Lie algebra $L_{\mathfrak{G}}$, see [KrZa24, Sect. 7.1], where
$L_{\mathfrak{G}}={\rm Span}_{\mathbb{K}}\{T_{e}Y\mid Y\subseteq\mathfrak{G}\,\,\text{a closed algebraic subset smooth in}\,\,e\}.$
Consider also the Zariski closure $\overline{G}$ in the ind-group $\operatorname{Aut}(X)$. This is a closed ind-subgroup with Lie algebra $\operatorname{Lie}(\overline{G})$. Notice that $L(G)\subseteq L_{G}$ and both are ideals in $\operatorname{Lie}\overline{G}$, see [KrZa24, Theorem 2.3.1 and Proposition 7.1.3].
Question 2.1 ([KrZa24, Question 1]).
Do we have $L(G)=L_{G}=\operatorname{Lie}(\overline{G})$?
The answer is affirmative provided $L$ is finite-dimensional.
Theorem 2.2 ([CD03, Theorem 1], [KrZa24, Theorem A]).
Assume that $L(G)$ is finite dimensional. Then $G$ is an algebraic subgroup of $\operatorname{Aut}(X)$ and $\operatorname{Lie}(G)=L_{G}=L(G)$.
2.2. Locally finite endomorphisms and Lie subalgebras
Definition 2.3.
An endomorphism $\lambda$ of a vector space $V$ is called locally finite if every $v\in V$ belongs to a finite-dimensional $\lambda$-invariant subspace of $V$.
Every locally finite endomorphism $\lambda$ has a uniquely defined additive Jordan decomposition $\lambda=\lambda_{s}+\lambda_{n}$ where $\lambda_{s}$ is semisimple, $\lambda_{n}$ locally nilpotent and $\lambda_{s}\circ\lambda_{n}=\lambda_{n}\circ\lambda_{s}$. Recall that a locally finite endomorphism $\lambda_{s}$ is semisimple if the restriction of $\lambda_{s}$ to any finite-dimensional $\lambda_{s}$-invariant subspace of $V$ can be diagonalized.
Definition 2.4.
A subspace $L\subseteq\operatorname{Vec}(X)=\operatorname{Der}(\mathcal{O}(X))$ is called locally finite if any $f\in\mathcal{O}(X)$ is contained in a finite dimensional $L$-invariant subspace of $\mathcal{O}(X)$.
Every element of a locally finite subspace is locally finite as an endomorphism of $\mathcal{O}(X)$ and thus admits a Jordan decomposition.
Lemma 2.5 ([KrZa24, Lemma 1.6.2]).
A locally finite subspace $L\subseteq\operatorname{Vec}(X)$ is finite dimensional.
Let $L\subseteq\operatorname{Vec}(X)$ be a Lie subalgebra generated by a family of locally finite Lie subalgebras $L_{i}\subseteq\operatorname{Vec}(X)$, $i\in I$. There is the following question, see [KrZa24, Question 2].
Question 2.6.
Is $L$ locally finite provided $L$ is finite dimensional?
According to Theorem 2.2 the answer is positive if $L_{i}=\operatorname{Lie}(G_{i})$, where the $G_{i}$ are algebraic subgroups of $\operatorname{Aut}(X)$. In particular, we have the following corollary.
Corollary 2.7 ([KrZa24, Corollary of Theorem E]).
Consider a family $\{\eta_{i}\mid i\in I\}$ of locally nilpotent vector fields on $X$ and let $L:=\langle\eta_{i}\mid i\in I\rangle_{\text{\tiny Lie}}\subseteq\operatorname{Vec}(X)$ be the Lie algebra generated by the $\eta_{i}$. If $L$ is finite dimensional, then $L$ is locally finite.
Recall that a locally nilpotent vector field $\eta$ on $X$ defines a one-dimensional algebraic subgroup $U=\exp(t\eta)\subseteq\operatorname{Aut}(X)$ isomorphic to $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$ such that $\operatorname{Lie}U={\mathbb{K}}\,\eta$.
A weak form of Question 2.6 is as follows.
Question 2.8.
Let $\xi,\eta\in\operatorname{Vec}(X)$ be locally finite. Assume that the Lie subalgebra $L=\langle\xi,\eta\rangle_{\text{\tiny Lie}}$ generated by $\xi$ and $\eta$ is finite dimensional. Does it follow that $L$ is locally finite?
By Corollary 2.7 the latter is true if $\xi$ and $\eta$ are locally nilpotent. Indeed, under the latter assumption we have by Theorem 2.2 $L=\operatorname{Lie}(G)$ where $G=\langle\exp(t\xi),\exp(t\eta)\rangle$ is an algebraic group.
3. Bijective morphisms of ind-groups
[…]
Definition 3.1 ([FuKr18, Definition 1.9.4]).
Let $\mathfrak{V}$ be an ind-variety, and let $x\in\mathfrak{V}$. We say that
  • $\mathfrak{V}$ is strongly smooth in $x$ if there is an open neighborhood of $x$ which has an admissible filtration consisting of smooth connected algebraic varieties.
  • $\mathfrak{V}$ is geometrically smooth in $x$ if there is an admissible filtration $\mathfrak{V}=\bigcup_{k}V_{d}$ such that $x\in V_{d}$ is a smooth point for all $d$.
Proposition 3.2 ([FuKr18, Propositions 1.8.5, 1.9.6 and 2.4.1]).
Assume that ${\mathbb{K}}$ is uncountable. Let $\varphi\colon\mathfrak{V}\to\mathfrak{W}$ be a bijective morphism of ind-varieties.
  • (a)
    Suppose that $\mathfrak{V}$ is connected and that $\mathfrak{W}$ has an admissible filtration $\mathfrak{W}=\bigcup_{k}W_{k}$ such that all $W_{k}$ are irreducible and normal. Then $\varphi$ is an isomorphism.
  • (b)
    Suppose that $\mathfrak{V}$ is curve-connected and $\mathfrak{W}$ is strongly smooth in a point $w\in\mathfrak{W}$. Then there is an open neighborhood $\mathfrak{V}^{\prime}$ of $\varphi^{-1}(w)$ in $\mathfrak{V}$ such that $\varphi|_{\mathfrak{V}^{\prime}}\colon\mathfrak{V}^{\prime}\to\mathfrak{W}$ is an open immersion. If $\mathfrak{W}$ is strongly smooth in every point, then $\varphi$ is an isomorphism.
Let now $\varphi\colon\mathfrak{G}\to\mathfrak{H}$ be a bijective homomorphism of ind-groups. Assume that $\mathfrak{G}$ is connected and $\mathfrak{H}$ is strongly smooth in $e$. Then $\varphi$ is an isomorphism.
Example 3.3 (see [FuKr18, Proposition 14.2.1]).
The automorphism group $\operatorname{Aut}({\mathbb{K}}\langle x,y\rangle)$ of the rank $2$ free associative algebra ${\mathbb{K}}\langle x,y\rangle$ has a natural structure of an ind-group. The homomorphism of abelianization ${\mathbb{K}}\langle x,y\rangle\to{\mathbb{K}}[x,y]$ induces a bijective (homo)morphism of ind-groups
$\varphi\colon\operatorname{Aut}({\mathbb{K}}\langle x,y\rangle)\to\operatorname{Aut}({\mathbb{K}}[x,y]),$
see [ML70], [Cze71-72] and [Coh85, Theorem 9.3]. The induced surjective morphism of Lie algebras
$\operatorname{Lie}(\operatorname{Aut}({\mathbb{K}}\langle x,y\rangle))\to\operatorname{Lie}(\operatorname{Aut}({\mathbb{K}}[x,y]))$
has a non-trivial kernel, see [BW00, last paragraph of sect. 11] and [FuKr18, Proposition 14.2.1]. In particular, $\varphi$ is not an isomorphism of ind-groups. In fact, there exists a unicuspidal curve $C$ in $\operatorname{Aut}({\mathbb{K}}\langle x,y\rangle)$ such that $\varphi(C)\subseteq\operatorname{Aut}({\mathbb{K}}[x,y])$ has a deeper cusp and so, is not isomorphic to $C$, see [FuKr18, Sect. 14.3].
Using Proposition 3.2 one arrives at the following conclusion.
Corollary 3.4 ([FuKr18, Corollary 14.1.2]).
Let ${\mathbb{K}}$ be uncountable. Then the ind-group $\mathfrak{G}=\operatorname{Aut}({\mathbb{K}}[x,y])$ is not strongly smooth in $\operatorname{id}$. Moreover, there is no admissible filtration $\mathfrak{G}=\cup_{d}\mathfrak{G}_{d}$ with irreducible and normal algebraic varieties $\mathfrak{G}_{d}$.
Remark 3.5.
It is not clear how the notions “strongly smooth” and “geometrically smooth” are related. There exists as well a notion of algebraic smoothness of a point in an ind-variety. It has the following advantage: the point $e\in\mathfrak{G}$ is algebraically smooth in every affine ind-group $\mathfrak{G}$, see [Sha81, Theorem 3] and [Kum02, Theorem 4.3.7]. A geometrically smooth point of an ind-variety is algebraically smooth, see [Kum02, Lemma 4.3.4].
There are several pathological examples. For instance, there is an ind-variety $\mathfrak{V}$ such that each point $v\in\mathfrak{V}$ is not algebraically smooth, see [Sha81, P. 188]. The ind-group $\mathfrak{G}=\operatorname{SL}(2,{\mathbb{K}}[t])$ has an admissible ascending filtration $\mathfrak{G}=\cup_{d}G_{d}$ such that the neutral element $e$ of $\mathfrak{G}$ is singular in every $G_{d}$, see [Sha04, Theorem 2] and [Kum02, Examples 4.3.5 and 4.3.8].
4. Nested ind-groups
Definition 4.1.
A subgroup $G$ of an ind-group $\mathfrak{G}$ is nested if $G$ is a union of an increasing sequence of algebraic subgroups $G_{i}$ of $\mathfrak{G}$, see [KPZ17]. A nested ind-group $\mathfrak{G}$ is called unipotent if all the $G_{i}$ are unipotent algebraic groups.
The last assertion of Proposition 3.2 can be applied to nested ind-groups. Indeed, such a group is strongly smooth, see [FuKr18, Example 2.4.2]. Notice that a connected nested ind-group is algebraically generated, and any closed ind-subgroup of a nested ind-group is nested.
Question 4.2 ([KrZa24, Question 4]).
Is it true that every connected nested ind-subgroup $G\subseteq\operatorname{Aut}(X)$ is closed in $\operatorname{Aut}(X)$?
The answer is affirmative, see [Per24, Theorem 6.6]. Furthermore, a subgroup of $\operatorname{Aut}(X)$ consisting of unipotent elements is nested if and only if it is closed in $\operatorname{Aut}(X)$, see [Per24, Theorem 8.11]. The proofs of these results exploit a new notion of a de Jonquières-like (dJ-like, for short) subgroup of $\operatorname{Aut}(X)$. The following analogue of the Lie–Kolchin theorem holds: every nested unipotent subgroup $U\subseteq\operatorname{Aut}(X)$ is contained in a dJ-like subgroup, see [Per24, Theorem 6.1]. Yet another ingredient in the proof of [Per24, Theorem 6.6] is the following version of the Levi decomposition for nested subgroups, see [KPZ17, Theorem 2.11 and Corollary 2.12].
Theorem 4.3.
[…]
There is the following conjectural criterion of nestedness.
Conjecture 4.4 (cf. [FuKr18, 9.4.3-9.4.6]).
Assume that every finite subset of an ind-group $\mathfrak{G}$ is contained in an algebraic group. Then $\mathfrak{G}$ is nested.
Recall the following definition.
Definition 4.5 ([FuKr18, Definition 1.13.2]).
A subset $S$ of an ind-group $\mathfrak{G}$ is called weakly closed if for every algebraic subset $U\subseteq\mathfrak{G}$ such that $U\subseteq S$ we have $\overline{U}\subseteq S$. The union $\cup_{U}\overline{U}$ for all $U$ as before is called the weak closure of $S$.
Notice that a countable union of closed algebraic subsets in $\mathfrak{G}$ is weakly closed, see [FuKr18, Proposition 1.13.6].
Question 4.6 ([FuKr18, Question 15.9.1]).
Let $g\in\operatorname{Aut}({\mathbb{A}}^{n})$. Are the following assertions equivalent?
  • $g$ is diagonalizable.
  • $g$ is semisimple.
  • The conjugacy class $C(g)$ is closed in $\operatorname{Aut}({\mathbb{A}}^{n})$.
  • The conjugacy class $C(g)$ is weakly closed in $\operatorname{Aut}({\mathbb{A}}^{n})$.
The next proposition contains a partial answer to Question 4.6.
Proposition 4.7 ([FuKr18, Corollary 15.9.8]).
Let $g\in\operatorname{Aut}({\mathbb{A}}^{n})$ be semisimple. Then $g$ is diagonalizable if and only if its conjugacy class $C(g)$ is weakly closed.
Question 4.8 ([FuKr18, Question 15.9.11]; cf. [FuKr18, Proposition 15.9.12]).
Let $g\in\operatorname{Aut}({\mathbb{A}}^{n})$ be locally finite. Does the (weak) closure of $C(g)$ contain the semisimple part of $g$?
5. Algebraic subgroups and algebraic elements
Definition 5.1.
An element $g$ of an ind-group $\mathfrak{G}$ is called algebraic if it is contained in an algebraic subgroup of $\mathfrak{G}$. It is called unipotent if it is contained in a $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup of $\mathfrak{G}$.
Clearly, $g$ is algebraic iff the closure $\overline{<g>}$ is an algebraic group. If $g$ is algebraic, then this abelian group has the form $(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{s}(\operatorname{{\mathbb{G}}_{\mathrm{a}}})^{t}F$ with $s\geq 0$, $t\in\{0,1\}$ and $F$ being a finite cyclic group. Every subgroup of this form in $\mathfrak{G}$ is equal to $\overline{<g>}$ for some $g\in\mathfrak{G}$.
It is easily seen that $g\in\operatorname{Aut}(X)$ is algebraic if and only if it defines a locally finite automorphism of $\mathcal{O}(X)$. In particular, any algebraic element $g$ has a unique multiplicative Jordan decomposition $g=g_{\mathrm{s}}g_{\mathrm{u}}$, where $g_{\mathrm{s}},g_{\mathrm{u}}\in\operatorname{Aut}(X)$ commute, $g_{\mathrm{s}}$ is semisimple and $g_{\mathrm{u}}$ is unipotent, see [FuKr18, Section 9.1].
Let $\operatorname{Aut}(X)_{\mathrm{alg}}\subseteq\operatorname{Aut}(X)$ stand for the subgroup generated by all algebraic elements, in other words, by all algebraic subgroups of $\operatorname{Aut}(X)$.
Theorem 5.2 ([BD15, Theorem 1.1]).
Assume that the field ${\mathbb{K}}$ is uncountable. Let $S$ be an affine surface given in ${\mathbb{A}}^{4}=\operatorname{Spec}({\mathbb{K}}[x,y,z,u])$ by the equations
$yu=xP(x),\quad vx=uQ(u),\quad yv=P(x)Q(u),$
where each of the polynomials $P,Q\in{\mathbb{K}}[w]$ has at least 2 distinct roots and $P(0)\neq 0$. Then the following hold.
  • The normal subgroup $\operatorname{Aut}(S)_{\mathrm{alg}}\subseteq\operatorname{Aut}(S)$ is not generated by a countable family of algebraic subgroups, and
  • the quotient $\operatorname{Aut}(S)/\operatorname{Aut}(S)_{\mathrm{alg}}$ contains a free group over an uncountable set of generators.
Conjecture 5.3.
Let $G=\langle G_{1},\ldots,G_{n}\rangle$ be a subgroup of $\operatorname{Aut}(X)$ generated by a finite collection of algebraic subgroups $G_{i}$. Then either $G$ is an algebraic group, or $G$ contains non-algebraic elements.
A weak version:
Conjecture 5.4.
Let $g,h\in\operatorname{Aut}(X)$ be algebraic elements. Then either $g,h$ are contained in an algebraic subgroup of $\operatorname{Aut}(X)$, or the group generated by $g$ and $h$ contains non-algebraic elements.
The following corollary of [LPU23, Theorem 5.1] confirms the latter conjecture in dimension two.
Theorem 5.5.
Let $X$ be an affine surface and $G$ be a finitely generated subgroup of $\operatorname{Aut}(X)$. If $G$ consists of algebraic elements, then $G$ is contained in an algebraic subgroup of $\operatorname{Aut}(X)$.
Cf. also Theorem 8.7 and Corollary 8.8 below.
Due to the next results (see [BPZ24, Theorem 1.1 and Corollaries 1.2-1.3]), the Burnside problem has a positive solution in $\operatorname{Aut}(X)$.
Theorem 5.6.
Let $k$ be any field of zero characteristic and $X$ be an affine variety over $k$. If $G$ is a torsion subgroup of $\operatorname{Aut}(X)$, then any finitely generated subgroup of $G$ is finite. If $X={\mathbb{A}}^{n}_{k}$, then $G$ contains an abelian normal subgroup of finite index.
The proof uses the following result of Bass and Lubotzky [BL83, Corollary 1.2].
Theorem 5.7.
Let $k$ be an arbitrary field and $X$ be a scheme of finite type over $k$. If $G\subseteq\operatorname{Aut}(X)$ is a finitely generated subgroup, then $G$ is residually finite, i.e. the intersection of all finite index subgroups of $G$ is trivial. If $X$ is flat over ${\mathbb{Z}}$, then $G$ is virtually torsion free, in particular, the orders of finite subgroups of $G$ are bounded above.
Example 5.8 (see [KrZa24, Theorem 6.4.2]).
Let $F=\langle g,h\rangle\subseteq\operatorname{Aut}({\mathbb{A}}^{2})$ be the subgroup generated by algebraic elements
$g\colon(x,y)\mapsto(x+y^{2},y)\quad\text{and}\quad h\colon(x,y)\mapsto(x,y+x^{2}).$
Then $F\cong{\mathbb{F}}_{2}$ is a free group of rank 2. It contains non-algebraic elements, for example, $gh$ is one of them. The closure $\mathcal{F}=\overline{F}$ is a free product $\mathcal{F}=\mathcal{J}^{+}*\mathcal{J}^{-}$ of two nested unipotent closed subgroups $\mathcal{J}^{\pm}$ of $\operatorname{Aut}({\mathbb{A}}^{2})$ isomorphic to the vector group $({\mathbb{K}}[t],+)$. Every algebraic subgroup of $\mathcal{F}$ is conjugate to a subgroup of $\mathcal{J}^{+}$ or $\mathcal{J}^{-}$.
Conjecture 5.9 ([FuKr18, Question 9.1.5]).
Assume that $\operatorname{Aut}^{0}(X)\neq\{\operatorname{id}_{X}\}$. Then $\operatorname{Aut}^{0}(X)$ contains a subgroup isomorphic to either $\mathbb{G}_{\mathrm{m}}$ or $\mathbb{G}_{\mathrm{a}}$, or in other words, $\operatorname{Aut}(X)$ contains a non-torsion algebraic element.
Due to Theorem 8.4 below, Conjecture 5.9 holds for affine surfaces. It holds also provided $\operatorname{Aut}(X)$ contains a nontrivial connected set of commuting elements. Indeed, we have the following results.
Theorem 5.10 ([CRX19, Theorem B]).
Let $Y$ be an irreducible algebraic subvariety of $\operatorname{Aut}(X)$ such that $Y$ contains the identity and any $g_{1},g_{2}\in Y$ commute. Then the subgroup of $\operatorname{Aut}(X)$ generated by $Y$ is an abelian connected algebraic subgroup of $\operatorname{Aut}(X)$.
Corollary 5.11.
Let $G$ be a connected abelian subgroup of $\operatorname{Aut}(X)$.
  • If $G$ is nontrivial, then it contains a subgroup isomorphic to either $\mathbb{G}_{\mathrm{m}}$ or $\mathbb{G}_{\mathrm{a}}$.
  • If $G$ is a closed ind-subgroup, then $G$ is nested.
As a consequence, an eventual counterexample to Conjecture 5.9 has the property that every commutative subgroup of $\operatorname{Aut}(X)$ is discrete, or, in other terms, every algebraic element of $\operatorname{Aut}(X)$ is a torsion element. See also [RvS21, Theorem A] for a description of maximal abelian subgroups of $\operatorname{Aut}(X)$ which consist of unipotent elements.
Question 5.12.
Assume that the group $\operatorname{Aut}(X)$ does not contain algebraic elements different from $\operatorname{id}_{X}$. Is it then discrete?
Notice that by Theorem 5.10, if $\operatorname{Aut}(X)$ has no algebraic element different from $\operatorname{id}_{X}$, then every abelian subgroup $H$ of $\operatorname{Aut}(X)$ has no torsion and the neutral component $H^{0}$ is trivial. Cf. also [Arz18a, Problem 20].
6. Nilpotency and solvability
Theorem 6.1 ([KrZa24, Theorem B]).
Let $G\subseteq\operatorname{Aut}(X)$ be an algebraically generated subgroup, see subsection 2.1. If $G$ is solvable, then the following hold.
  • (a)
    $G=U\rtimes T$ where $T$ is an algebraic torus and $U=R_{\mathrm{u}}(G)$ is a nested unipotent group.
  • (b)
    $G$ is unipotent if and only if the generating subgroups are.
  • (c)
    If, in addition, $G$ is generated by a finite family of connected algebraic groups, then $G$ is a connected algebraic group.
Theorem 6.2 ([KrZa24, Theorem C]).
A nested unipotent subgroup $U\subseteq\operatorname{Aut}(X)$ is solvable of derived length $\leq\max\{\dim Ux\mid x\in X\}\leq\dim X$.
Remark 6.3.
A unipotent algebraic group is nilpotent. By contrast, a nested unipotent group $G=\cup_{i}G_{i}$ is not necessarily nilpotent, since the nilpotency class of the unipotent algebraic subgroups $G_{i}$ might not be bounded.
For example, consider the de Jonquières subgroup ${\rm Jonq}_{n}\subseteq\operatorname{Aut}({\mathbb{A}}^{n})$ of triangular automorphisms of the form
$\varphi=(a_{1}x_{1}+f_{1},a_{2}x_{2}+f_{2}(x_{1}),\ldots,a_{n}x_{n}+f_{n}(x_{1},\ldots,x_{n-1}))$
where $a_{i}\in{\mathbb{K}}^{*}$ and $f_{i}\in{\mathbb{K}}[x_{1},\ldots,x_{i-1}]$. The unipotent radical $\mathcal{J}_{n}=R_{\mathrm{u}}({\rm Jonq}_{n})$ is a closed nested unipotent subgroup of $\operatorname{Aut}({\mathbb{A}}^{n})$. For $n\geq 2$ its Lie algebra $\operatorname{Lie}(\mathcal{J}_{n})$ is not nilpotent. However, the Lie algebra of a nilpotent ind-group is nilpotent, see [KrZa24, Lemma 5.1.4(3)].
Indeed, let for instance $n=2$ and
$L_{d}=\langle\partial/\partial x_{1},x_{1}^{d}\partial/\partial x_{2}\rangle_{Lie}\subseteq\operatorname{Lie}(\mathcal{J}_{2}).$
It is easily seen that the $d$th member $(L_{d})_{d}$ of the lower central series of $L_{d}$ does not vanish, and so $(\operatorname{Lie}(\mathcal{J}_{2}))_{d}\neq 0$ for any $d\geq 1$, cf. [KrZa24, Remark 5.3.3].
Question 6.4.
Let $\mathfrak{G}$ be a connected ind-group. Is it true that $\mathfrak{G}$ is solvable (resp., nilpotent) if $\operatorname{Lie}\mathfrak{G}$ is?
Definition 6.5.
Let us say that a subgroup $G\subseteq\operatorname{Aut}(X)$ is a-generated if $G=\langle Y\rangle$ where $Y\subseteq\operatorname{Aut}(X)$ is an irreducible algebraic subset containing $\operatorname{id}_{X}$. Replacing $Y$ by $Y\cdot Y^{-1}$ we may assume that $Y$ is symmetric, that is, $Y=Y^{-1}$.
Conjecture 6.6.
If $G$ is a-generated and nilpotent (resp., solvable) then $G$ is algebraic.
In the case of a nilpotent (resp., a solvable) subgroup $G$ one can try to proceed by induction on the nilpotency class of $G$ (resp., on the derived length of $G$), where the case $n(G)=1$ (i.e. $G$ abelian) is settled by Theorem 5.10. In the general case, the induction works provided the following conjecture is true.
Conjecture 6.7.
If $G$ is a-generated and nilpotent (resp., solvable) then any member of its lower central series (resp., derived series) is a-generated as well.
Let $z(G)$ stands for the center of $G$. If $G/z(G)$ is abelian, then Conjecture 6.6 holds if the following is true:
Conjecture 6.8.
Consider an admissible filtration $\operatorname{Aut}^{0}(X)=\bigcup A_{n}$ by irreducible affine subvarieties $A_{n}$. Let $G$ be a (nilpotent, resp., solvable) subgroup of $\operatorname{Aut}^{0}(X)$ such that $G=\langle G\cap A_{n}\rangle$ for some $n$. Then $z(G)=\langle z(G)\cap A_{m}\rangle$ for some $m$ (resp., $(G,G)=\langle(G,G)\cap A_{m}\rangle$ for some $m$).
The following question concerns a modified version of stable triangulation, see Section 9.8.
Question 6.9 ([KrZa24, Sect. 5.2, Question 6]).
Let $G\subseteq\operatorname{Aut}(X)$ be a solvable or nilpotent connected subgroup. Does there exists a closed embedding $X\hookrightarrow{\mathbb{A}}^{n}$ such that $G$ extends to a subgroup of the de Jonquières group ${\rm Jonq}_{n}$? Is this true if $G$ is algebraic or nested?
Remark 6.10.
The answer to the latter question is affirmative in a particular setup where $X={\mathbb{A}}^{k}$ and $G\subseteq\operatorname{Aut}({\mathbb{A}}^{k})$ is a unipotent algebraic group, see [KrZa24, Remark 1.4.2(3)]. For instance, the Nagata-Bass $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup of $\operatorname{Aut}({\mathbb{A}}^{3})$ can be triangulated in ${\mathbb{A}}^{4}$ in the above sense, being non-stably triangulable in the usual sense, cf. Example 9.37.
Denote by $\operatorname{LND}(X)\subseteq\operatorname{Vec}(X)$ the set of all locally nilpotent derivations of $\mathcal{O}(X)$. For a vector space $V$ over ${\mathbb{K}}$ we let $\operatorname{LND}(V)$ be the set of locally nilpotent endomorphisms of $V$.
Definition 6.11 ([Dai19]).
A subset $Z$ of endomorphisms of a vector space $V$ is called uniformly locally nilpotent, or ULN for short, if for any $v\in V$ there is $n\in{\mathbb{N}}$ such that $a_{1}\cdots a_{n}(v)=0$ whatever are $a_{1},…,a_{n}\in Z$.
Clearly, if $Z$ is ULN then $Z\subseteq\operatorname{LND}(V)$ and ${\rm span}_{{\mathbb{K}}}\,(Z)\subseteq\operatorname{End}_{\mathbb{K}}(V)$ is ULN too.
Let $A\subseteq\operatorname{Aut}(X)$ be an irreducible algebraic subvariety which contains $\operatorname{id}_{X}$, and let $\langle A\rangle$ be the subgroup of $\operatorname{Aut}(X)$ generated by $A$. The notation $L_{A}\subseteq\operatorname{Vec}(X)$ has the same meaning as in subsection 2.1. Given an ind-subgroup $\mathfrak{G}$, the adjoint representation $\operatorname{Ad}(\mathfrak{G})\colon\mathfrak{G}\to\operatorname{GL}(\operatorname{Lie}\mathfrak{G})$ is the tangent representation at the fixed point $\operatorname{id}_{X}\in\mathfrak{G}$ of the action of $\mathfrak{G}$ on itself by inner automorphisms ${\rm Int}(g)\colon h\mapsto g\cdot h\cdot g^{-1}$, see [FuKr18, Section 7.5].
Question 6.12.
Do the following equivalences hold?
  1. (1)
    $L_{A}\subseteq\operatorname{LND}(X)\Leftrightarrow\operatorname{Ad}(A)\subseteq\operatorname{End}_{Lie}(\operatorname{Vec}(X))$ is ULN $\Leftrightarrow\langle A\rangle$ is a unipotent algebraic group;
  2. (2)
    $\langle A\rangle$ is metabelian $\Leftrightarrow\langle\operatorname{Ad}(A)\rangle_{Lie}$ is commutative;
  3. (3)
    $\langle A\rangle$ is solvable $\Leftrightarrow\langle\operatorname{Ad}(A)\rangle_{Lie}$ is solvable;
  4. (4)
    $\langle A\rangle$ is nilpotent $\Leftrightarrow\langle\operatorname{Ad}(A)\rangle_{Lie}$ is nilpotent.
7. $\operatorname{Aut}$-quasihomogeneous and flexible affine varieties
7.1. $\operatorname{Aut}$-quasihomogeneous affine varieties
Definition 7.1.
We say that a (not necessarily affine) variety $X$ is $\operatorname{Aut}$-homogeneous (resp. $\operatorname{Aut}$-quasihomogeneous) if $\operatorname{Aut}(X)$ acts transitively on $X$ (resp. acts on $X$ with an open orbit).
The following characterizations of $\operatorname{Aut}$-quasihomogeneity for smooth affine surfaces are due to Gizatullin [Giz71b]. This was extended to normal affine surfaces by Dubouloz [Dub04].
Theorem 7.2.
Let $X$ be a normal affine surface $X$ non-isomorphic to ${\mathbb{A}}^{1}\times({\mathbb{A}}^{1}\setminus\{0\})$. Then the following are equivalent:
  • $X$ is $\operatorname{Aut}$-quasihomogeneous;
  • there are two effective $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-actions on $X$ with distinct general orbits;
  • $X\cong\bar{X}\setminus D$ where $\bar{X}$ is a normal projective surface and $D\subseteq\bar{X}$ is a reduced divisor with simple normal crossings contained in the smooth locus ${\rm reg}(\bar{X})$ of $\bar{X}$, which consists of a chain of smooth rational curves.
Notice that many, but not all, smooth Gizatullin surfaces are $\operatorname{Aut}$-homogeneous. See Gizatullin [Giz71a] and Popov [Pop73] for a list of normal affine surfaces that are (quasi)homogeneous spaces of algebraic groups; see also [KPZ17, Theorem 4.8].
In higher dimensions there is a classification of $\operatorname{Aut}$-homogeneous affine toric varieties and complete toric varieties, see [Arz18b]; cf. also the recent preprint [ArZa23].
7.2. Flexible affine varieties
Flexible affine varieties were introduced in [AFK+13a], see also [KZ99] and [AKZ12]. For overviews on the subject and examples we address [Arz23], [AFK+13b], and [CPPZ21, Section 7].
Definition 7.3.
Let $X$ be a quasiaffine variety. We say that a smooth point $x\in X$ is flexible if the tangent space $T_{x}X$ is spanned by the tangent vectors to the orbits $Ux$ of $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups $U$ of $\operatorname{Aut}(X)$. The variety $X$ is called flexible if every smooth point $x\in X$ is flexible. Clearly, $X$ is flexible if at least one smooth point $x\in X$ is and the group $\operatorname{Aut}(X)$ acts transitively on the smooth locus ${\rm reg}(X)$.
We let $\operatorname{\mathfrak{U}}(X)\subseteq\operatorname{Aut}^{0}(X)$ be the subgroup generated by all $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups of $\operatorname{Aut}(X)$, in other words, by all unipotent elements of $\operatorname{Aut}(X)$  [^1] The subgroup $\operatorname{\mathfrak{U}}(X)$ is denoted by ${\rm SAut}(X)$ in [AFK+13a]. We prefer the notation $\operatorname{\mathfrak{U}}(X)$ in order to keep ${\rm SAut}({\mathbb{A}}^{n})$ for the subgroup of $\operatorname{Aut}({\mathbb{A}}^{n})$ of volume-preserving automorphisms.. We also let $\operatorname{Aut}_{\rm alg}(X)$ be the subgroup generated by all algebraic elements of $\operatorname{Aut}(X)$. Notice that $\operatorname{Aut}_{\rm alg}(X)$ is not necessarily connected, and every algebraic element $a\in\operatorname{Aut}_{\rm alg}(X)\setminus\operatorname{Aut}^{0}(X)$ is a torsion element.
The following theorem (conjectured in [AKZ12, Sect. 4.2]) was proven in [AFK+13a] for affine varieties and extended in [APS14, Theorem 2] and [FKZ16, Theorem 2.12] to quasiaffine varieties.
Theorem 7.4.
For a quasiaffine variety $X$ of dimension at least 2, the following are equivalent:
  • $X$ is flexible;
  • the group $\operatorname{\mathfrak{U}}(X)$ acts transitively on the smooth locus ${\rm reg}(X)$;
  • the group $\operatorname{\mathfrak{U}}(X)$ acts infinitely transitively on ${\rm reg}(X)$.
The flexibility survives when removing a subvariety of codimension $\geq 2$.
Theorem 7.5 ([FKZ16, Theorem 0.1]).
Let $X$ be a flexible smooth quasiaffine variety of dimension $\geq 2$, and let $Y\subseteq X$ be a closed subset of codimension $\geq 2$. Then $X\setminus Y$ is flexible.
As an application of Theorem 7.4, let us mention the following two interpolation results. The first follows immediately from [AFK+13a, Theorem 4.14 and Remark 4.15].
Theorem 7.6.
Let $X$ be a flexible affine variety of dimension $n\geq 2$. Then for every $m\geq 0$ and every finite subset $Z\subseteq{\rm reg}(X)$ there exists an automorphism $g\in\operatorname{\mathfrak{U}}(X)$ with prescribed $m$-jets at the points $p\in Z$ provided each jet fixes its center $p$.
Corollary 7.7.
Under the assumptions of Theorem 7.6 there exists an ${\mathbb{A}}^{1}$-orbit of a $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-action on $X$ which passes through every point $p\in Z$ and interpolates a given smooth curve jet at $p$.
See also [AFK+13a, Theorem 1.15] for an analogue of Kleiman’s Transversality theorem for flexible varieties.
The first nontrivial examples of flexible varieties appeared in [Giz71a] and [Giz71b]; these are $\operatorname{Aut}$-homogeneous Gizatullin affine surfaces, see subsection 7.1. See [GD75], [GD77], [KPZ17] and references therein for systematic studies on the automorphism groups of Gizatullin surfaces.
Let us mention two interesting examples of flexible affine varieties of the form $X_{n,d}={\mathbb{P}}^{n}\setminus H_{d}$, where $H_{d}$ is a hypersurface of degree $d$ in ${\mathbb{P}}^{n}$ with $(n,d)\in\{(26,3),\,(55,4)\}$, see [Giz18, Proposition 3.7 and Theorem 4.3]. These are homogeneous spaces of simple linear groups of type $E_{6}$ and $E_{7}$, respectively. The cubic form that defines $H_{3}\subseteq{\mathbb{P}}^{26}$ was first written by Camille Jordan in 1870. It is related to the configuration of $27$ lines on a smooth cubic surface in ${\mathbb{P}}^{3}$. The quartic form that defines $H_{4}\subseteq{\mathbb{P}}^{55}$ is related to the $56$ exceptional curves on a Del Pezzo surface of degree 2, and to the 28 bitangent lines of a plane quartic curve.
7.3. Flexibility of toric varieties
Let ${\mathbb{T}}=(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{n}$ stand for an algebraic $n$-torus. Recall that a ${\mathbb{T}}$-variety $X$ of dimension $n$ is a toric variety if ${\mathbb{T}}$ acts effectively on $X$ with an open orbit. One says that a toric variety $X$ has a torus factor if $X\cong Y\times(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{k}$ for some toric variety $Y$ and some $k>0$, or equivalently, if there is a nonconstant invertible regular function on $X$.
Any toric affine variety $X$ of dimension $n\geq 2$ with no torus factor can be defined by a strictly convex polyhedral cone in the integer lattice ${\mathbb{Z}}^{n}$. For instance, the affine plane ${\mathbb{A}}^{2}$ with the standard diagonal torus action is defined by the first quadrant ({$({\mathbb{N}}\cup\{0\})^{2}$ of ${\mathbb{Z}}^{2}$.
Definition 7.8.
Let $X$ be a toric variety. A Demazure root subgroup of $\operatorname{Aut}(X)$ is a $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup $U\subseteq\operatorname{Aut}(X)$ which is normalized by the acting torus ${\mathbb{T}}$.
Such subgroups, or rather the associated locally nilpotent derivations of $\mathcal{O}(X)$, correspond to certain lattice points of ${\mathbb{Z}}^{n}$ called Demazure roots. In particular, if $X={\mathbb{A}}^{2}$ then the Demazure roots are the lattice points $(-1,k)$ and $(l,-1)$ with $k,l\geq 0$. They correspond to the locally nilpotent derivations $y^{k}\partial/\partial x$ and $x^{l}\partial/\partial y$, respectively. It is known that any normal toric affine variety of dimension $\geq 2$ with no torus factor is flexible, see [AKZ12]. The following theorem says more.
Theorem 7.9 ([AKZ19, Theorem 1.3]).
Let $X$ be a toric affine variety of dimension at least 2. If $X$ has no torus factor and is smooth in codimension $2$, then one can find a finite collection of Demazure root subgroups such that the group generated by these acts infinitely transitively on the smooth locus ${\rm reg}(X)$.
8. Rigid affine varieties
Definition 8.1.
An affine variety $X$ is called rigid if it admits no effective $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-action.
Example 8.2.
Conjecturally, a Pham–Brieskorn hypersurface $X\subseteq{\mathbb{A}}^{n+1}$, $n\geq 2$, defined by
$x_{0}^{a_{0}}+x_{1}^{a_{1}}+\cdots+x_{n}^{a_{n}}=0\quad\text{where}\quad 2\leq a_{0}\leq a_{1}\leq\cdots\leq a_{n}$
is rigid if and only if $a_{1}\geq 3$, see [CPPZ21, Conjecture 1.22]. This is indeed the case for $n=2$ [KZ00, Lemma 4] and for $n=3$ [ChDu23, Main theorem]. For example, the Fermat threefold hypersurface
$x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3}=0$
in ${\mathbb{A}}^{4}$ is rigid; see [CPPZ21, Theorem 1.13] for a far more general result. See also [EGS23] and references therein for a description of rigid trinomial affine varieties.
For the following conjecture see e.g. [PZ22, Conjectures 1.0.1 and 1.0.2]; cf. also Conjecture 5.9 and Question 5.12.
Conjecture 8.3.
Let $X$ be a rigid affine variety. Then $\operatorname{Aut}^{0}(X)$ is an algebraic torus (of rank $\leq\dim(X)$).
Conjecture 8.3 is confirmed for rigid toric varieties and varieties with a torus action of complexity one, see [BG23]. Due to the following theorem, this conjecture also holds in dimension $2$.
Theorem 8.4 ([PZ22, Theorem 1.3(1)]).
Assume that the base field ${\mathbb{K}}$ is uncountable. Let $X$ be a normal affine surface over ${\mathbb{K}}$. Then $\operatorname{Aut}^{0}(X)$ is an algebraic group if and only if $X$ is rigid, if and only if $\operatorname{Aut}^{0}(X)$ is an algebraic torus (of rank $\leq 2$).
We have the following geometric criterion of rigidity of a normal affine surface.
Theorem 8.5 ([PZ22, Theorem 1.3(2)]).
Assume that the base field ${\mathbb{K}}$ is uncountable. Then a normal affine surface $X$ over ${\mathbb{K}}$ is rigid if and only if $X$ admits a relatively minimal completion $\bar{X}$ by a reduced effective normal crossing divisor $D$ supported on ${\rm reg}(\bar{X})$ such that any extremal chain of the dual graph $\Gamma(D)$ which consists of rational components $C_{1},\ldots,C_{k}$ satisfies $C_{i}^{2}\leq-2$ for $i=1,\ldots,k$.
The dual graph $\Gamma(D)$ is defined as follows: its vertices correspond to components of $D$ and the edges correspond to the pairs of components that intersect. A chain $L$ in $\Gamma(D)$ is called extremal if all vertices of $L$ are of valence $\leq 2$ in $\Gamma(D)$ and at least one vertex is of valence 1.
Example 8.6 ([PZ22, Example 4.3.6]).
Consider the affine surface $X=\mathbb{P}^{2}\setminus{\rm supp}(D)$ where $D$ is a reduced effective divisor on $\mathbb{P}^{2}$ with only nodes as singularities. Then $X$ is not rigid if and only if $\deg(D)\leq 2$, that is, $D$ is either a projective line of a (possibly, singular) conic.
By definition, a nested ind-group consists of algebraic elements. The following theorem provides a partial converse.
Theorem 8.7 ([PR24, Theorem 1.1], cf. also [PR23]).
For an affine variety $X$ the following conditions are equivalent:
  • $\operatorname{Aut}^{0}(X)$ consists of algebraic elements;
  • $\operatorname{Aut}^{0}(X)$ is a closed nested ind-subgroup of $\operatorname{Aut}(X)$;
  • $\operatorname{Aut}^{0}(X)=\operatorname{\mathfrak{U}}(X)\rtimes T$ where $T$ is a maximal torus of $\operatorname{Aut}(X)$ and $\operatorname{\mathfrak{U}}(X)=R_{u}(\operatorname{Aut}(X))$ is abelian and consists of all unipotent elements of $\operatorname{Aut}(X)$.
Corollary 8.8 ([PR24, Corollary 4.3]).
Let $X$ be a rigid affine variety. If $\operatorname{Aut}^{0}(X)$ consists of algebraic elements, then $\operatorname{Aut}^{0}(X)$ is an algebraic torus of dimension $\leq\dim(X)$.
The proof exploits the following result of independent interest.
Theorem 8.9 ([PR24, Theorem 3.1]).
Let $\mathfrak{G}$ be a connected ind-group and $\mathfrak{H}\subseteq\mathfrak{G}$ be a closed connected nested ind-subgroup. Assume that any $g\in\mathfrak{G}$ is a torsion element modulo $\mathfrak{H}$, that is, $g^{d}\in\mathfrak{H}$ for some $d=d(g)>0$. Then $\mathfrak{H}=\mathfrak{G}$.
The following theorem gives a necessary condition for rigidity. For the proof see [AG17, Theorem 2.1]; cf. also [FZ05b, Theorem 3.3] in the case of affine surfaces.
Theorem 8.10.
Suppose that $\operatorname{Aut}(X)$ contains two non-commuting $\operatorname{{\mathbb{G}}_{\mathrm{m}}}$-subgroups. Then $\operatorname{Aut}(X)$ contains a $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup.
This theorem implies (a) below; see [PR24, Proposition 5.1 and Remark 5.2] for the proof of (b).
Theorem 8.11.
Let $X$ be a rigid affine variety. Then the following hold.
  • (a)
    There exists an algebraic torus ${\mathbb{T}}\cong(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{k}$ in $\operatorname{Aut}(X)$ which contains any other algebraic subtorus of $\operatorname{Aut}(X)$. In particular, ${\mathbb{T}}$ is a normal subgroup of $\operatorname{Aut}(X)$.
  • (b)
    ${\mathbb{T}}$ is a central subgroup of $\operatorname{Aut}^{0}(X)$ and the only maximal connected abelian ind-subgroup of $\operatorname{Aut}^{0}(X)$.
The next theorem confirms Conjecture 8.3 in a particular setting of rigid varieties with a torus action. For statement (a) see [AG17, Example 2.3 and Remark 2.4] and [AKZ12, Sect. 2]; (b) is proven in [BG23, Theorem 6.4]. Recall that the complexity of an algebraic group action is the codimension of general orbits.
Theorem 8.12.
  • (a)
    Let $X$ be a normal toric variety. If $X$ is rigid, then $X\cong{\mathbb{T}}$ is a torus and $\operatorname{Aut}^{0}(X)\cong{\mathbb{T}}$ is the torus acting naturally on itself.
  • (b)
    Let $X$ be a normal rational affine variety with a torus action of complexity one and a finitely generated divisor class group. Suppose that $X$ is rigid and every invertible regular function on $X$ is constant. Then $\operatorname{Aut}^{0}(X)$ is an algebraic torus.
Concerning the conjugacy of tori in $\operatorname{Aut}(X)$ we have the following general results, see e.g. [FuKr18, Proposition 10.5.4] for (a) and [Dem82], [Gub98] and [BH03] for (b).
Theorem 8.13.
  • (a)
    Let $T\subseteq\operatorname{Aut}(X)$ be an algebraic torus. Then $\dim(T)\leq\dim(X)$.
  • (b)
    Assume that $X$ is a normal toric variety. Then all algebraic tori $T\subseteq\operatorname{Aut}(X)$ of dimension $\dim(X)$ are conjugate. Furthermore, any algebraic torus of dimension $\dim(X)-1$ in $\operatorname{Aut}(X)$ is contained in an algebraic torus of dimension $\dim(X)$.
However, for a non-toric affine variety $X$ the group $\operatorname{Aut}(X)$ can contain several pairwise non-conjugate maximal algebraic tori.
Example 8.14.
  1. (See [Dani89, Theorem 2].) Let $D$ be the Danielewski surface $\{xy-z^{2}+z=0\}$ in ${\mathbb{A}}^{3}_{{\mathbb{C}}}$ and let $X=D\times{\mathbb{A}}^{1}$. Then $\operatorname{Aut}(X)$ contains an infinite sequence of pairwise non-conjugate algebraic $2$-tori and no $3$-torus. Furthermore, let $X$ be the underlying affine variety of $\operatorname{SL}(2,{\mathbb{C}})\times{\mathbb{A}}^{1}_{{\mathbb{C}}}$. Then the number of conjugate classes of maximal tori in $\operatorname{Aut}(X)$ is infinite [Dani89, P. 7].
2. There exist examples of $\operatorname{Aut}$-homogeneous smooth non-toric affine surfaces $X$ such that $\operatorname{Aut}(X)$ contains an algebraic family of pairwise non-conjugate algebraic $1$-tori depending on 1 or 2 parameters, see [FKZ11, Theorem 1.0.1] and also [KPZ17, Theorem 5.5]. Cf. also Example 9.21 below of Danilov-Gizatullin surfaces, where the number of conjugate classes of $1$-tori is finite.
Remark 8.15 (see [KML97] and [DMJP10]).
Recall that the Koras-Russel threefold $X$ given in ${\mathbb{A}}^{4}_{{\mathbb{C}}}$ by equation $x+x^{2}y+z^{2}+t^{3}=0$ is diffeomorphic to ${\mathbb{A}}^{3}_{{\mathbb{C}}}$, but non-isomorphic to ${\mathbb{A}}^{3}_{{\mathbb{C}}}$. It is neither flexible, nor rigid. The group $\operatorname{Aut}(X)$ is infinite dimensional, acts on $X$ with an open orbit and a unique fixed point (the origin). It coincides with the stabilizer of $X$ in $\operatorname{Aut}({\mathbb{A}}^{4}_{{\mathbb{C}}})$. This group is naturally isomorphic to the subgroup of $\operatorname{Aut}({\mathbb{A}}^{3}_{{\mathbb{C}}})$ (where ${\mathbb{A}}^{3}_{{\mathbb{C}}}=\operatorname{Spec}{\mathbb{C}}[x,z,t]$) of all automorphisms which leave invariant the ideals $(x)$ and $(x^{2},x+z^{2}+t^{3})$.
9. Chapters on $\operatorname{Aut}({\mathbb{A}}^{n})$
9.1. Groups acting infinitely transitively on ${\mathbb{A}}^{n}$
Recall that the root subgroups of $\operatorname{Aut}({\mathbb{A}}^{n})$ are the $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups normalized by the standard $n$-torus, see Definition 7.8.
Theorem 9.1 ([And19], [AKZ19]).
For any $n\geq 2$ one can find three $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups (resp., $n+2$ root subgroups) of $\operatorname{Aut}({\mathbb{A}}^{n})$ which generate a subgroup acting infinitely transitively on ${\mathbb{A}}^{n}$.
9.2. Tame subgroups
Notice that the de Jonquières subgroup ${\rm Jonq}_{n}\subseteq\operatorname{Aut}({\mathbb{A}}^{n})$ (see Remark 6.3) is generated by the standard torus ${\mathbb{T}}$ of $\operatorname{GL}(n,{\mathbb{K}})$ and the unipotent radical $\mathcal{J}_{n}=R_{\mathrm{u}}({\rm Jonq}_{n})$. In turn, $\mathcal{J}_{n}$ is generated by the triangular root subgroups
$\exp(tx_{1}^{m_{1}}\cdots x_{i-1}^{m_{i-1}}\partial_{i}),\quad i=1,\ldots,n\quad\text{where}\quad\partial_{i}=\partial/\partial x_{i}.$
The Lie algebra
$\operatorname{Lie}(\mathcal{J}_{n})={\mathbb{K}}\partial_{1}\oplus{\mathbb{K}}[x_{1}]\partial_{2}\oplus\cdots\oplus{\mathbb{K}}[x_{1},\ldots,x_{n-1}]\partial_{n}$
consists of locally nilpotent derivations. This is a maximal Lie subalgebra in $\operatorname{Vec}({\mathbb{A}}^{n})$ consisting of locally nilpotent derivations, see [Sku21, Theorems 1 and 7]; cf. also [BPZ24, Theorems 1.4 and 1.5] and the references therein for closely related results. For ${\mathbb{K}}={\mathbb{C}}$, ${\rm Jonq}_{n}$ is a Borel subgroup of $\operatorname{Aut}({\mathbb{A}}_{{\mathbb{C}}}^{n})$, that is, a maximal (connected) solvable subgroup, see [FuPo18, Corollary 1.2].
Definition 9.2.
An automorphism $\varphi\in\operatorname{Aut}({\mathbb{A}}^{n})$ is said to be tame if it is a composition of affine and triangular automorphisms; otherwise it is called wild. The tame automorphisms form the tame subgroup $\mathop{\rm Tame}_{n}=\langle\operatorname{Aff}_{n},{\rm Jonq}_{n}\rangle$ of $\operatorname{Aut}({\mathbb{A}}^{n})$.
For the following classical theorem see [Jun42], [vdK53]; cf. also [Kam75, Theorem 2] and [Miy24, Theorem 2.3.8].
Theorem 9.3 (Jung-van der Kulk).
Let $k$ be an arbitrary field. Then Aut($\operatorname{Aut}({\mathbb{A}}^{2}_{k})=\mathop{\rm Tame}_{2}$. Moreover, Aut($\operatorname{Aut}({\mathbb{A}}^{2}_{k})$ is a free product $\operatorname{Aff}_{2}*_{C}{\rm Jonq}_{2}$ amalgamated over $C=\operatorname{Aff}_{2}\cap{\rm Jonq}_{2}$.
Using this amalgam, Danilov [Dan74] showed that the subgroup ${\rm SAut}({\mathbb{A}}^{2}_{\mathbb{K}})=\{f\in\operatorname{Aut}({\mathbb{A}}^{2}_{\mathbb{K}})|\det(f)=1\}$ of unimodular automorphisms is not a simple group.
The things are different for Aut($\operatorname{Aut}({\mathbb{A}}^{3}_{\mathbb{K}})$. Nagata considered in [Nag72, Section 2.1] the following automorphism of ${\mathbb{A}}^{3}_{\mathbb{K}}=\operatorname{Spec}({\mathbb{K}}[x,y,z])$:
$\Phi\colon(x,y,z)\mapsto(x+\sigma z,y+2\sigma x+\sigma^{2}z,z)\quad\text{where}\quad\sigma=x^{2}-yz.$
We have
$\Phi=\exp(\sigma\partial)\quad\text{where}\quad\partial=z\partial/\partial x+2x\partial/\partial y.$
Indeed, $\partial(\sigma)=0$ and so, $\Phi^{*}(\sigma)=\sigma$. Since also $\Phi^{*}(z)=z$ we have $\Phi\in\operatorname{Aut}_{{\mathbb{K}}[z]}{\mathbb{K}}[z][x,y]$. Nagata showed that $\Phi$ is not generated by the affine and triangular automorphisms that fix $z$, and suggested that $\Phi$ is wild, i.e., $\Phi\notin\mathop{\rm Tame}_{3}$, see [Nag72, Conjecture 3.1]. The following theorem due to Shestakov and Umirbaev confirms Nagata’s Conjecture.
Theorem 9.4 ([SU04a, SU04b]; see also [vdEKC21, Chapter 1]).
The Nagata automorphism $\Phi$ is wild. So, $\mathop{\rm Tame}_{3}$ is a proper subgroup of $\operatorname{Aut}({\mathbb{A}}^{3})$.
Notice that $\mathop{\rm Tame}_{3}$ is not a closed ind-subgroup of $\operatorname{Aut}({\mathbb{A}}^{3})$, see [EP15]. It is unknown whether $\mathop{\rm Tame}_{3}$ is dense in $\operatorname{Aut}({\mathbb{A}}^{3})$.
Example 9.5 ([FuKr18, Sect. 17.3]).
Letting ${\mathbb{A}}^{3}=\operatorname{Spec}{\mathbb{K}}[x,y,z]$ consider the closed connected ind-subgroup $\mathfrak{G}\subseteq\operatorname{Aut}({\mathbb{A}}^{3})$ consisting of the automorphisms which send $z$ to $z$. Then $\mathfrak{G}^{t}=\mathfrak{G}\cap\mathop{\rm Tame}_{3}$ is a proper closed ind-subgroup of $\mathfrak{G}$. Nevertheless, their Lie algebras coincide: $\operatorname{Lie}(\mathfrak{G}^{t})=\operatorname{Lie}(\mathfrak{G})$, contrary to the claim of [Sha81, Theorem 1].
Question 9.6.
Do we have $L_{\mathfrak{G}}=\operatorname{Lie}\mathfrak{G}$ and $L_{\mathfrak{G}^{t}}=\operatorname{Lie}\mathfrak{G}^{t}$? Is it true that $L_{\mathfrak{G}}=L_{\mathfrak{G}^{t}}$?
Let us mention also the following result.
Theorem 9.7 ([KrSt13, Main Theorem]).
Every automorphism of the group $\operatorname{Aut}({\mathbb{A}}^{n}_{\mathbb{C}})$ is inner up to field automorphisms when restricted to the tame subgroup $\mathop{\rm Tame}_{n}$.
Remarks 9.8.
1. For $n=2$ the assertion was established in [Dés06]. Since $\mathop{\rm Tame}_{2}=\operatorname{Aut}({\mathbb{A}}^{2}_{\mathbb{C}})$, see Theorem 9.3, it follows that every automorphism of $\operatorname{Aut}({\mathbb{A}}^{2}_{\mathbb{C}})$ is inner up to a field automorphism. The latter does not hold, in general, for toric affine surfaces. For example, this fails for the quotient of ${\mathbb{A}}^{2}$ by the cyclic group generated by $g\colon(x,y)\mapsto(\xi^{e}x,\xi y)$, where $\xi$ is a $d$th primitive root of unity, $1<e<d$, $(e,d)=1$ and $e^{2}\not\equiv 1\mod d$, see [LRU23, Remark 5.13].
2. Any automorphism of $\operatorname{Aut}({\mathbb{A}}^{n})$ as an ind-group is inner, see [BKYE16]. The same holds for the automorphisms of the subgroup ${\rm SAut}({\mathbb{A}}^{n})$ of volume-preserving automorphisms of ${\mathbb{A}}^{n}$, see [Kra17]. See also [Sta13], [UZ21, Theorem 1.3] and the literature therein for some generalizations of Theorem 9.7.
3. A tame subgroup was defined for some other automorphism groups, for instance,
  • for $\operatorname{Aut}(Q_{3})$ where $Q_{3}$ is a smooth affine quadric threefold realized as the underlying variety of $\operatorname{SL}(2,{\mathbb{C}})$, see [LV13];
  • for $\operatorname{Aut}(A_{n})$ where $A_{n}={\mathbb{K}}\langle x_{1},\ldots,x_{n}\rangle$ is the free associative algebra in $n\geq 2$ variables, see [BKY12b];
  • for $\operatorname{Aut}(X)$, where $X$ is a toric affine variety, see [AG10, Definition 5.3].
A theorem due to Makar-Limanov and Czerniakiewicz says that $\operatorname{Aut}(A_{2})=\mathop{\rm Tame}(\operatorname{Aut}(A_{2}))$, see [ML70], [Cze71-72] and [Coh85, Theorem 4.1]. At the same time, $\mathop{\rm Tame}(\operatorname{Aut}(Q_{3}))$ and $\mathop{\rm Tame}(\operatorname{Aut}(A_{3}))$ are proper subgroups of $\operatorname{Aut}(Q_{3})$ and $\operatorname{Aut}(A_{3})$, respectively, see [AG10, Theorem 6.1] and [LV13, Section 5.1] for the former and [Umi07, Corollary 2.1] for the latter. The role of the Nagata automorphism for $\operatorname{Aut}(A_{3})$ is played by the Anick automorphism
$(x,y,z)\mapsto(x+z\sigma,y+x\sigma,z)\quad\text{where}\quad\sigma=xz-zy,$
which is wild. See also [SZ24] for a wild (but stably tame) automorphism of Anick type of a $3$-generated free Poisson algebra, which induces a tame automorphism of the 3-generated polynomial algebra.
Likewise, consider the quadratic cone $S_{3}\subseteq{\mathbb{A}}^{4}$ given by equation $x_{1}x_{4}-x_{2}x_{3}=0$. There exists an automorphism $\tau$ of ${\mathbb{A}}^{4}$ which restricts to $S_{3}$ and lifts to the Cox ring ${\mathbb{K}}[x_{1},x_{2},x_{3},x_{4}]$ of $S_{3}$ yielding the Anick automorphism of ${\mathbb{A}}^{4}$:
$(x_{1},x_{2},x_{3},x_{4})\mapsto(x_{1},x_{2}+x_{1}\delta,x_{3},x_{4}+x_{3}\delta)\quad\text{where}\quad\delta=x_{1}x_{4}-x_{2}x_{3}.$
The restriction $\tau|_{S_{3}}$ is a wild automorphism of $S_{3}$, see [AG10, Theorem 6.1].
The following theorem relates the automorphism group of ${\mathbb{A}}^{n}$ with the automorphism groups of various Lie algebras of vector fields on ${\mathbb{A}}^{n}$. Below $\operatorname{Vec}^{0}({\mathbb{A}}^{n})$ stands for the Lie algebra of vector fields with zero divergence.
Theorem 9.9.
There are isomorphisms
$\operatorname{Aut}({\mathbb{A}}^{n})\cong\operatorname{Aut}_{\rm Lie}(\operatorname{Vec}({\mathbb{A}}^{n}))\cong\operatorname{Aut}_{\rm Lie}(\operatorname{Vec}^{\mathrm{c}}({\mathbb{A}}^{n}))\cong\operatorname{Aut}_{\rm Lie}(\operatorname{Vec}^{0}({\mathbb{A}}^{n}))$
induced by the natural action of the group $\operatorname{Aut}({\mathbb{A}}^{n})$ on the Lie algebra of vector fields $\operatorname{Vec}({\mathbb{A}}^{n})$.
The first isomorphism was obtained in [Rud86, Corollary], [Kul93, Sec. 4] and [Bav17a] along different approaches; in [Kul93] it appeared in relation with the Jacobian Conjecture. The other isomorphisms were established independently in [Bav17b] and [KrRe17]. See [Bav14] for the automorphism group of the Lie algebra of triangular derivations.
See also [CL06, Théorème A] for a description of the lattices in simple real Lie groups that can be realized as subgroups of $\operatorname{Aut}({\mathbb{A}}^{2}_{{\mathbb{C}}})$. Furthermore, see [AGL24] and the references therein for studies on (finite dimensional) Lie algebras of derivations of polynomial rings.
9.3. Tamizable automorphisms
Definition 9.10 ([MP09, Definition 4.1]).
An automorphism $\varphi\in\operatorname{Aut}({\mathbb{A}}^{n})$ is called tamizable (resp., linearizable, triangulable) if it is conjugate in $\operatorname{Aut}({\mathbb{A}}^{n})$ to a tame (resp. linear, triangular) automorphism.
Question 9.11 ([MP09]).
Is it true that any $\varphi\in\operatorname{Aut}({\mathbb{A}}^{n})$ is tamizable? In particular, is the Nagata automorphism $\Phi$ tamizable?
Example 9.12.
It is known that the Nagata automorphism $\Phi$ is not conjugate to a triangular one, see [Bas84]. Nevertheless, it is shifted linearizable. More precisely, $2{\rm id}\cdot\Phi$ is conjugate to $2{\rm id}$ in $\operatorname{Aut}({\mathbb{A}}^{3})$, see [MP09, Theorem 3.3]; cf. also [FuKr18, Lemma 15.11.1] for a general result of this type.
As follows from the Jung-van der Kulk Theorem 9.3, $\operatorname{Aut}({\mathbb{A}}^{2})$ is algebraically generated, see subsection 2.1 for the definition. It is known that $\operatorname{Aut}({\mathbb{A}}^{n})$ is connected for any $n$, see [Sha81, Lemma 4]. The following natural question arises:
Question 9.13.
Is $\operatorname{Aut}({\mathbb{A}}^{n})$ algebraically generated for $n\geq 3$?
The locally finite automorphisms of $\operatorname{Aut}({\mathbb{A}}^{n})$ were studied in [FuMa07]. It is known that the subgroup $\langle{\rm LF}_{n}\rangle$ of $\operatorname{Aut}({\mathbb{A}}^{n})$ generated by all locally finite automorphisms is normal, see [Zyg11]. By the Jung-van der Kulk Theorem 9.3 we have $\langle{\rm LF}_{2}\rangle=\operatorname{Aut}({\mathbb{A}}^{2})$. For $n=3$, $\langle{\rm LF}_{3}\rangle$ contains the Nagata automorphism. However, it is unknown whether $\operatorname{Aut}({\mathbb{A}}^{3})=\langle{\rm LF}_{3}\rangle$ ([Zyg11]; cf. also subsection 9.5).
9.4. Stable tameness
Definition 9.14.
A wild automorphism $\varphi\in\operatorname{Aut}({\mathbb{A}}^{n})$ is called stably tame if $\varphi\times{\rm\operatorname{id}}_{{\mathbb{A}}^{\kappa}}\in\operatorname{Aut}({\mathbb{A}}^{n+\kappa})$ is tame for some $\kappa\geq 1$. The minimal number $\kappa$ of additional variables will be called excessive dimension.
Theorem 9.15 ([Smi89]).
The Nagata automorphism $\Phi$ is stably tame with excessive dimension $\kappa=1$.
By [Spo07] the same is true for the Anick automorphism of ${\mathbb{A}}^{4}$, cf. Remark 9.8.3. More generally, the following holds.
Theorem 9.16 ([BvdEW12, Corollary 4.9]).
Let as before $\mathfrak{G}\subseteq\operatorname{Aut}({\mathbb{A}}^{3})$ be the subgroup of automorphisms fixing $z$. Then any element $\varphi\in\mathfrak{G}$ is stably tame with excessive dimension $\kappa\leq 3$.
A similar result is also known for $\operatorname{Aut}(A_{3})$ with $\kappa=1$, see [BKY12b, Theorem 1.1].
9.5. Cotame automorphisms
Definition 9.17 (Cotame automorphisms; cf. [EL19]).
One says that $h\in\operatorname{Aut}({\mathbb{A}}^{n})\setminus{\operatorname{Aff}}_{n}$ is cotame if $\langle{\operatorname{Aff}}_{n},h\rangle\supseteq{\mathop{\rm Tame}}_{n}$ and topologically cotame if $\overline{\langle{\operatorname{Aff}}_{n},h\rangle}\supseteq{\mathop{\rm Tame}}_{n}$.
Theorem 9.18 ([Bod02], [Bod05, 1.4 and Theorem 1.8], [EL19]).
For $n\geq 3$ any triangular $h\in{\rm Jonq}_{n}\setminus\operatorname{Aff}_{n}$ is cotame, while there is no triangular cotame $h\in\operatorname{Aut}({\mathbb{A}}^{2})\setminus\operatorname{Aff}_{2}$.
The following theorem of Edo extends and refines the earlier results of Bodnarchuk ([Bod01, Thm. 3]) and Furter ([Fur15, Thm. D]); cf. also [FuPo18, Theorem 1.4].
Theorem 9.19 ([Edo18, Thm. 1.2]).
For $n\geq 2$ any element $h\in\operatorname{Aut}({\mathbb{A}}^{n})\setminus\operatorname{Aff}_{n}$ is topologically cotame.
9.6. Amalgamation
By Jung-van der Kulk’s Theorem 9.3, $\operatorname{Aut}({\mathbb{A}}^{2})$ is an amalgam of two subgroups along their intersection. A similar decomposition exists for the group $\operatorname{GL}(2,{\mathbb{K}}[x])$, see [Nag59], and for the group $\operatorname{Aut}^{0}(A_{2})$ of augmentation-preserving automorphisms of the free associative algebra $A_{2}$, see [SY98, Theorem 1.1].
It is unknown, however, whether a similar result holds for $\operatorname{Aut}({\mathbb{A}}^{n})$ when $n\geq 3$. For the tame subgroup $\mathop{\rm Tame}_{3}\subseteq\operatorname{Aut}({\mathbb{A}}_{3})$ we have the following result.
Theorem 9.20 ([Wri15], see also [Wri17]).
$\mathop{\rm Tame}_{3}$ can be realized as a generalized amalgamated product of three subgroups amalgamated along pairwise intersections.
See [ANU20, Theorem 1] and the references therein for analogues of the Jung-van der Kulk’s Theorem for the automorphism group of the free Lie algebra in three variables and the subgroup of tame automorphisms. Similar results are also known to hold for certain $\operatorname{Aut}$-homogeneous affine surfaces among the so-called Danilov-Gizatullin surfaces.
Example 9.21 (Danilov-Gizatullin surfaces).
These are smooth affine surfaces $X_{d}=F_{n}\setminus S_{d}$ where $\pi_{n}\colon F_{n}\to{\mathbb{P}}^{1}$ is the $n$th (smooth projective) Hirzebruch ruled surface and the curve $S_{d}\subseteq F_{n}$ is an ample section of $\pi_{n}$ with $S_{d}^{2}=d>n$. Two such affine surfaces are isomorphic if and only if they share the same invariant $d$, see [GD77, Theorem 5.8.1]. The group $\operatorname{Aut}(X_{d})$ has exactly $\lfloor d/2\rfloor$ conjugacy classes of maximal tori of rank 1, see [FKZ07, Prop. 5.15]. The group $\operatorname{Aut}(X_{d})$ is an amalgam of 2 nested ind-subgroups if $d=3$ and of 3 nested ind-subgroups if $d=4,5$, see [GD77, Sections 5-8]. Assume that the base field ${\mathbb{K}}$ is uncountable. Then $\operatorname{Aut}(X_{6})$ is not an amalgam of two nested ind-groups. For $d\geq 7$, $\operatorname{Aut}(X_{d})$ is not an amalgam of any countable number of nested ind-groups, see [KPZ17, Corollary 5.6].
One can find similar results for other $\operatorname{Aut}$-quasihomogeneous affine surfaces in [GD75, GD77], see also [KPZ17, Sect. 5].
9.7. Linearization of reductive group actions
Let $G$ be a reductive linear algebraic group acting effectively on ${\mathbb{A}}^{n}$ with a fixed point $x$. Then the tangent representation of $G_{x}$ on $T_{x}X$ is faithful. Luna’s Étale Slice Theorem [Lun73] guarantees a local linearization of the $G_{x}$-action near $x$. This justifies an expectation that any action of a reductive group on ${\mathbb{A}}^{n}$ can be linearized, that is, is conjugate in $\operatorname{Aut}({\mathbb{A}}^{n})$ to a linear action; see e.g. [Kam79, Conjecture 3.1] and [vdE00, Section 9.4]. However, the latter expectation turns out to be wrong as we will see below in 9.7.
The first results in this context are due to Białynicki-Birula who proved the following in [Bia66-67].
Theorem 9.22.
A faithful action of a torus of dimension $\geq n-1$ on ${\mathbb{A}}^{n}$ is linearizable. In particular, all maximal tori in $\operatorname{Aut}({\mathbb{A}}^{n})$ are pairwise conjugate.
The latter fails for affine surfaces and affine threefolds, in general, see Examples 8.14 and 9.21.
It then took a long time to settle the next open problem, namely the linearizability of a $\mathbb{G}_{\mathrm{m}}$-action on ${\mathbb{A}}^{3}_{{\mathbb{C}}}$. This was done in a series of fundamental papers of Koras and Russell, see e. g. [KoRu97], completed finally in [KKMLR97].
Theorem 9.23.
Any action of $\operatorname{{\mathbb{G}}_{\mathrm{m}}}$ on ${\mathbb{A}}^{3}_{\mathbb{C}}$ is linearizable.
The proof uses a classification of certain Koras-Russell affine threefolds diffeomorphic to $\mathbb{R}^{6}$ and non-isomorphic to ${\mathbb{A}}^{3}_{\mathbb{C}}$, see [KML97].
Certain actions of codimension 2 tori can be linearized also in higher dimensions, see [KamRu82] for (a) and [KoRu89] for (b).
Theorem 9.24.
Assume that the algebraic torus $T=(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{n-2}$ acts effectively on ${\mathbb{A}}^{n}$. Then this action can be linearized provided one of the following holds:
  1. (a)
    The action is ”unmixed,” that is, no nontrivial character, together with its inverse, occur as weights of semi-invariant polynomials;
  2. (b)
    the fixed point set of $T$ has positive dimension.
Next we look at the case of the affine plane ${\mathbb{A}}^{2}$. Using the amalgamated product structure of $\operatorname{Aut}({\mathbb{A}}^{2})$ and a famous result of Serre’s [Ser03, I.4.3.8] on subgroups of bounded length in an amalgamated product, one gets the following theorem.
Theorem 9.25 ([Kam79, Theorem 4.3]).
An algebraic subgroup $G$ of $\operatorname{Aut}({\mathbb{A}}^{2})$ is conjugate to a subgroup of either $\operatorname{Aff}_{2}$ or ${\rm Jonq}_{2}$. In particular, every reductive group action on ${\mathbb{A}}^{2}$ is linearizable.
In particular, any action of a finite group on ${\mathbb{A}}^{2}$ is conjugate to a linear action. This was first proven by geometric means by T. Igarashi (1975, unpublished), see also [Fur83, Theorem 2] and [Miy24, Theorem 2.4.6].
For actions of reductive groups we have the following consequence of Luna’s Étale Slice Theorem, see [Lun73, III. Corollaire 2]; cf. also [BH85, Corollary 10.7].
Theorem 9.26.
Let a reductive algebraic group $G$ act on an affine variety $X$. Assume that every $G$-invariant function is constant and that $G$ has a smooth fixed point in $X$. Then $X$ is $G$-equivariantly isomorphic to a representation of $G$.
Applied to the actions on ${\mathbb{A}}^{n}$, this gives the following corollary.
Corollary 9.27 ([KrPo85, Proposition 5.1]).
If a reductive group acts on ${\mathbb{A}}^{n}$ with a dense orbit, then the action is linearizable.
For semisimple groups we have the following results, see [KrPo85] and [KrRu14] for $n=3$ and [Pan84-86] for $n=4$.
Theorem 9.28.
For $n\leq 4$, any action of a semisimple algebraic group on ${\mathbb{A}}^{n}$ is linearizable.
In the paper [KrSc92] Kraft and Schwarz study the linearization problem for reductive group actions on smooth acyclic affine varieties $X$ assuming that the quotient $X/\!/G$ has dimension 1. It turns out that there is always a fixed point $x_{0}\in X$ ([KrSc92, Theorem 1]) so that the tangent representation $V:=T_{x_{0}}X$ is the candidate for the linearization. One of the main results is the following ([KrSc92, Theorem 5]).
Theorem 9.29.
$X$ is $G$-isomorphic to $V:=T_{x_{0}}X$ in the following cases:
  1. (1)
    $V$ is a semifree $G$-module;
  2. (2)
    $G$ is a torus;
  3. (3)
    $\dim V^{G^{0}}=1$;
  4. (4)
    $\dim X\leq 3$;
  5. (5)
    $G^{0}$ is semisimple;
  6. (6)
    $V$ is selfdual as a $G^{0}$-module.
As a corollary one gets the next result, cf. [KrRu14, Theorem A].
Corollary 9.30.
Every faithful action of a non-finite reductive group on ${\mathbb{A}}^{3}$ is linearizable.
As for linearizable finite group actions, Kraft and Procesi suggested a conjectural linearization procedure that uses the Reynolds operator. This linearization procedure works indeed in the case of involutions of ${\mathbb{A}}^{n}$ acting with a fixed point, provided the associated Taylor decomposition of the Reynolds average is of bounded length, see [Jur90].
Counterexamples
The first examples of non-linearizable actions of connected reductive groups on ${\mathbb{A}}^{n}$ are due to Schwarz [Sch89]. These are an ${\rm O}_{2}({\mathbb{C}})$-action on ${\mathbb{A}}_{{\mathbb{C}}}^{4}$ and an $\operatorname{SL}_{2}({\mathbb{C}})$-action on ${\mathbb{A}}_{{\mathbb{C}}}^{7}$. In a sense, the non-linearizability is a general phenomenon.
Theorem 9.31 ([Kno91]).
For any non-commutative connected reductive algebraic group $G$ there exists a non-linearizable $G$-action on some affine space ${\mathbb{A}}^{n}$.
Many more examples are given in [KrSc92, VII.5. Examples].
All known examples of non-linearizable actions on ${\mathbb{A}}^{n}$ are obtained from non-trivial algebraic $G$-vector bundles over $G$-modules, following an idea from [BH87], see also [Kra89], [KrSc92], [MP91] and [MMP91]. Indeed, by the Quillen-Suslin theorem, such a bundle is trivial when ignoring the $G$-action. Hence, its total space is isomorphic to some ${\mathbb{A}}^{n}$. The induced $G$-action on ${\mathbb{A}}^{n}$ is non-linearizable due to the fact that our bundle is non-trivial in the category of $G$-vector bundles.
However, this method does not work for an abelian reductive group $G$. Indeed, in this case every $G$-vector bundle over a $G$-module is trivial, see [MMP96]. The following questions remain open, see e.g. [KrSc95] and [Mas03]:
Question 9.32.
Is there a non-linearizable action of an abelian group on ${\mathbb{A}}^{n}$, for instance, of a cyclic group ${\mathbb{Z}}/p{\mathbb{Z}}$ on ${\mathbb{A}}^{n}$, $n\geq 3$ and of $\operatorname{{\mathbb{G}}_{\mathrm{m}}}$ on ${\mathbb{A}}^{n}$, $n\geq 4$?
Notice that there is an example of Asanuma of a non-linearizable action of $\operatorname{{\mathbb{G}}_{\mathrm{m}}}(\mathbb{R})$ on ${\mathbb{A}}^{n}_{\mathbb{R}}$, see [Asa99].
There are several examples of non-linearizable actions of non-abelian finite groups on ${\mathbb{A}}^{n}$, for instance, of the dihedral groups $D_{10}$, $D_{14}$ and of of the symmetric group $S_{3}$ on ${\mathbb{A}}^{4}$, see [MJ94, Theorem 5.1] and [FM15]; cf. also [MP95], [MMP96], [Mas03]. It seems that no example of this kind is known for a finite group action on ${\mathbb{A}}^{3}$, see e.g. [FM15]. See also [PeRa86] for the Lefschetz fixed-point formula in the context of finite group actions on affine varieties.
The linearization problem is closely related to the Abhyankar–Sathaye problem on linearization of embeddings ${\mathbb{A}}^{k}\hookrightarrow{\mathbb{A}}^{n}$ and on variables in polynomial rings. We send the reader e.g. to [DG15], [Kal09], [KrRu14], [vdEMV07] and the literature therein for further reading.
9.8. Solvable and unipotent ind-subgroups: triangulation
We start with the following result.
Proposition 9.33 ([FuKr18, Proposition 10.5.1]).
Let $U\subseteq\operatorname{Aut}(X)$ be a commutative unipotent algebraic subgroup of dimension $n$. Assume that ${\rm Cent}_{\operatorname{Aut}(X)}(U)=U$. Then for any $x\in X$ the orbit map $U\to X$, $u\mapsto ux$, is an isomorphism. In particular, $X$ is isomorphic to ${\mathbb{A}}^{n}$.
The unipotent elements of the de Jonquières group ${\rm Jonq}_{n}$ form a solvable subgroup $\mathcal{J}_{n}$ of derived length $n$, see [FuPo18, Lemma 3.2]. This subgroup is not nilpotent for $n>1$; see Remark 6.3 above.
Definition 9.34.
We say that a subgroup $G\subseteq\operatorname{Aut}({\mathbb{A}}^{n})$ is triangulable if it is conjugate in $\operatorname{Aut}({\mathbb{A}}^{n})$ to a subgroup of ${\rm Jonq}_{n}$. It is said to be stably triangulable if $G\times\operatorname{id}_{{\mathbb{A}}^{k}}$ is triangulable in $\operatorname{Aut}({\mathbb{A}}^{n+k})$ for some $k\geq 0$.
By [Ren68] any unipotent subgroup of $\operatorname{Aut}({\mathbb{A}}^{2})$ is triangulable, cf. Theorem 9.25. This is not any longer true in higher dimensions. However, the next theorem provides a large class of triangulable unipotent subgroups in any dimension.
Theorem 9.35 ([KrZa24, Theorem D]).
Let $U\subseteq\operatorname{Aut}({\mathbb{A}}^{n})$ be a nested unipotent subgroup. If $U$ acts transitively on ${\mathbb{A}}^{n}$, then $U$ is triangulable.
A similar result holds for Lie subalgebras of vector fields.
Theorem 9.36 ([Sku21, Theorem 6]).
Let $L\subseteq\operatorname{Lie}(\operatorname{Aut}({\mathbb{A}}^{n}))$ be a Lie subalgebra that consists of locally nilpotent derivations. Assume that $\bigcap_{\partial\in L}\ker(\partial)={\mathbb{K}}$. Then $L$ is ${\rm Ad}$-conjugate to a subalgebra of the Lie algebra $\operatorname{Lie}({\rm Jonq}_{n})$ of triangular derivations.
The transitivity assumption in Theorem 9.35 is essential, as the following example shows.
Example 9.37.
Consider the Nagata automorphism $\Phi=\exp(\sigma\partial)$, see subsection 9.2. It is contained in the $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup $U=\exp(t\sigma\partial)$ of $\operatorname{Aut}({\mathbb{A}}^{3})$. According to Bass [Bas84], $U$ is not triangulable. Indeed, the fixed points set $({\mathbb{A}}^{3})^{\operatorname{{\mathbb{G}}_{\mathrm{a}}}}$ is a hypersurface with an isolated singularity. However, the fixed point set of a triangulable $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup is isomorphic to a cylinder $X\times{\mathbb{A}}^{1}$ which cannot have an isolated singularity. Using the same idea one can construct non-triangulable $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups of any $\operatorname{Aut}({\mathbb{A}}^{n})$, $n\geq 3$, see [Pop87].
According to Freudenburg [Fre17, Lemma 3.36], the Nagata-Bass $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup $U$ is not stably triangulable in $\operatorname{Aut}({\mathbb{A}}^{4})$. In fact, no $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup of Bass-Popov type is stably triangulable, whatever is the number of additional variables.
Proposition 9.38 ([KrZa24, Proposition 1.4.1]).
Consider a $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup $U\subseteq\operatorname{Aut}({\mathbb{A}}^{n})$. Assume that the fixed point set $({\mathbb{A}}^{n})^{U}$ is a hypersurface with an isolated singularity. Then $U$ is not stably triangulable.
There is, however, the following positive result.
Theorem 9.39 ([Kal04]).
Assume that ${\mathbb{K}}={\mathbb{C}}$, and let a $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup $U\subseteq\operatorname{Aut}({\mathbb{A}}^{3})$ acts freely on ${\mathbb{A}}^{3}$. Then $U$ is conjugate in $\operatorname{Aut}({\mathbb{A}}^{3})$ to a subgroup of translations. In particular, $U$ is triangulable.
An analogue of the former assertion fails in ${\mathbb{A}}^{4}$, see [Win90]. For further results in this direction, see e.g. the survey article [Kal09] and [GMM21] for $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-actions on affine varieties, [Fre17] and [vdEKC21] for $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-actions on affine spaces and Hilbert’s 14th problem. See also [Pop17] and the references therein for (stably) birationally triangulable and non-triangulable actions of unipotent groups.
10. Tits’ type alternative for automorphism groups
A group $G$ is called virtually solvable resp., virtually nilpotent, etc. if $G$ is a finite extension of a solvable resp. nilpotent, etc. group. The classical Tits alternative says that any linear group $G$, that is a subgroup of $\operatorname{GL}(n,{\mathbb{K}})$, is either virtually solvable, or contains a free group ${\mathbb{F}}_{2}$ on two generators. For finitely generated subgroups, the Tits alternative holds without any restriction on the base field.
There are several similar statements for automorphism groups. For instance, an analogue of the Tits alternative holds for $\operatorname{Aut}({\mathbb{A}}^{2}_{{\mathbb{C}}})$, see [Lam01, Theorem 2.4] or [Lam24, Proposition 20.17].
Notice that Aut($\operatorname{Aut}({\mathbb{A}}^{2}_{{\mathbb{K}}})$ for an infinite field ${\mathbb{K}}$ is not isomorphic to a linear group over any extension of ${\mathbb{K}}$, and the same holds for its de Jonqières subgroup ${\rm Jonq}_{2}$, see [Cor17, Proposition 2.3 and Remark 2.4] and also [Mat23]. However, by [Mat23, Theorem A] the codimension 6 normal subgroup
$G=\{\varphi\in\operatorname{Aut}({\mathbb{A}}^{2}_{{\mathbb{K}}})\,|\,\varphi(0)=0,\,\,\,d\varphi(0)=\operatorname{id}\}$
of Aut($\operatorname{Aut}({\mathbb{A}}^{2}_{{\mathbb{K}}})$ is a linear group over a suitable field extension.
The unipotent triangular subgroup $\mathcal{J}_{n}$ of Aut($\operatorname{Aut}({\mathbb{A}}^{n}_{{\mathbb{K}}})$ over an arbitrary field ${\mathbb{K}}$ is a semidirect product of abelian groups, see [BNS12, Theorem 1]. For $n\geq 3$ this group is nonlinear (it cannot be represented by matrices over any field), see [Sos07, Theorem 2]. The same holds for the tame subgroup $\mathop{\rm Tame}_{n}=\mathop{\rm Tame}(\operatorname{Aut}({\mathbb{A}}^{n}_{{\mathbb{K}}}))$ provided ${\rm char}({\mathbb{K}})=0$ and $n\geq 4$, see [RCS04]. Similar results are also true for the automorphism groups of free associative algebras, see ibid.
For the proof of the following theorem see [Can11] and [Ure21] for (a), [BFL14, Theorem C] for (b) and [LP22, Main Theorem] for (c). In Theorems 10.1 and 10.2 we assume that the base field ${\mathbb{K}}$ is algebraically closed of characteristic zero.
Theorem 10.1.
The Tits alternative holds for the following groups:
  • (a)
    the birational automorphism group of any compact complex Kähler surface;
  • (b)
    $\mathop{\rm Tame}(\operatorname{Aut}(\operatorname{SL}(2,{\mathbb{K}})))$ (see Remark 9.8.3);
  • (c)
    $\mathop{\rm Tame}_{3}=\mathop{\rm Tame}(\operatorname{Aut}({\mathbb{A}}^{3}_{{\mathbb{K}}}))$.
In the case of algebraically generated groups of automorphisms, we have the following Tits’ type alternatives.
Theorem 10.2 ([AZ22, Theorem 1.1] and [AZ23, Theorem 6.1]).
  • (a)
    Let $X$ be an affine algebraic surface over ${\mathbb{K}}$ and $G\subseteq\operatorname{Aut}(X)$ be a subgroup generated by $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups $U_{1},\ldots,U_{k}$. Then either $G$ contains ${\mathbb{F}}_{2}$ or $G$ is a metabelian unipotent algebraic group.
  • (b)
    Let $X$ be a toric affine variety over ${\mathbb{K}}$ and $G\subseteq\operatorname{Aut}(X)$ be a subgroup generated by Demazure root subgroups $U_{1},\ldots,U_{k}$, see Definition 7.8. Then either $G$ contains ${\mathbb{F}}_{2}$ or $G$ is a unipotent algebraic group.
Similar results also hold for subgroups of $\mathop{\rm Tame}(\operatorname{Aut}(\operatorname{SL}(2,{\mathbb{K}})))$ and $\mathop{\rm Tame}(\operatorname{Aut}({\mathbb{A}}^{3}_{{\mathbb{K}}}))$, see [AZ23, Corollary 6.4].
Acknowledgments. We would like to thank Ivan Arzhantsev, Alexander Perepechko and Andriy Regeta for their comments and references. Our thanks also go to the anonymous referee for useful remarks.